if you are traveling 45 miles per hour how much distance do you cover in 3 seconds, to the nearest foot?
Answer by mananth(16949)

45 miles per hour how much distance do you cover in 3 seconds,
...
1 mph = 1.466 ft/second
45 mph = 45*1.466
= 66 ft/second
in 3 seconds it will travel = 66*3= 198 feet.
...
m.ananth@hotmail.ca